Assume perfect competition:

Price: $38

Cost: TC=20Q+.04Q2

TC=20Q+.04Q2

Solve for the profit-maximizing Quantity produced by an individual firm in the short run.
Answer #1

TC = 20Q + 0.04Q2

MC = 20 + 0.08Q

P = 38

The profit maximization condition is:

MC = P

20 + 0.08Q = 38

0.08Q = 38 - 20 = 18

Q = 18 / 0.08 = 225

Thus, the profit-maximizing Quantity produced by an individual firm is 225.